OpenStudy (anonymous):

PLEASE HEEELP!! find an equation of a circle that passes through (-1, 0), (0, 2) and (0, -1)

OpenStudy (anonymous):

This requires some work, but recall the general form of the equation of a circle: \[\Large f(x,y)=(x-u)^2+(y-v)^2=r^2 \] Where \(u,v\) are the midpoints of the circle and \(r\) is the radius of the circle. So you have three unknowns, but you also have three conditions in form of given points. Do you know what that means for your problem?

OpenStudy (anonymous):

look at this link,its solved in 3 different ways http://www.qc.edu.hk/math/Advanced%20Level/circle%20given%203%20points.htm